OpenStudy (anonymous):

If f(x)=2sinx^2x+sinx+1, find the value of f(π/6) 1.3/2 2. 2 3.5/2 4.5+- radical 3/2

OpenStudy (anonymous):

Plug in pi/6 and solve.

OpenStudy (anonymous):

it can be converted into degree by multiplying by 180 degree / pi

OpenStudy (anonymous):

so pi / 6 * 180 / pi = 30 degree now put 30 degree and solve it by putting values

OpenStudy (anonymous):

f(30) = 2sin^2 30degree + sin 30 degree + 1 = 1/2+1/2 +1 = 2